# WaterSeeker: Efficient Detection of Watermarked Segments in Large Documents

### v1

@@ Abstract
Watermarking algorithms for large language models (LLMs) have attained high accuracy in detecting LLM-generated text. However, existing methods primarily focus on distinguishing fully watermarked text from non-watermarked text, overlooking real-world scenarios where LLMs generate only small sections within large documents. In this scenario, balancing time complexity and detection performance poses significant challenges. This paper presents WaterSeeker, a novel approach to efficiently detect and locate watermarked segments amid extensive natural text. It first applies an efficient anomaly extraction method to preliminarily locate suspicious watermarked regions. Following this, it conducts a local traversal and performs full-text detection for more precise verification. Theoretical analysis and experimental results demonstrate that WaterSeeker achieves a superior balance between detection accuracy and computational efficiency. Moreover, WaterSeeker’s localization ability supports the development of interpretable AI detection systems. This work pioneers a new direction in watermarked segment detection, facilitating more reliable AI-generated content identification.
@@ 1 Introduction
As the quality of text generated by large language models (LLMs) improves, these models address various practical challenges while also raising concerns such as misinformation (Liu et al., 2023c; Chen & Shu, 2024) and copyright infringement (Rillig et al., 2023). To tackle these issues, LLM watermarking technology has emerged, embedding specific information (watermarks) during text generation for accurate detection through specialized algorithms. Existing watermarking methods mainly focus on distinguishing fully watermarked text from non-watermarked content (Kirchenbauer et al., 2023; Zhao et al., 2024; Liu et al., 2024a; Lu et al., 2024; Aaronson & Kirchner, 2022; Kuditipudi et al., 2024). However, in real-world scenarios, LLMs often generate only brief segments within longer documents, making it essential to detect these watermarked sections among large amounts of non-watermarked text.
Most previous algorithms relied on computing statistics across the entire document for detection, which we refer to as full-text detection methods. For instance, some approaches categorize vocabulary into green and red lists, applying a bias towards the selection of green tokens during generation and using metrics like the z-score to measure their proportion throughout the entire text (Kirchenbauer et al., 2023; Zhao et al., 2024; Liu et al., 2024a). Other methods utilize pseudo-random numbers to guide token sampling, calculating correlations across the entire document during detection (Aaronson & Kirchner, 2022; Christ et al., 2024; Kuditipudi et al., 2024). Ultimately, the presence of a watermark is determined by whether these statistics exceed a specified threshold.
@@ 2 Related Work
Watermarking for large language models (LLMs) embeds subtle, algorithmically detectable signals in outputs to identify LLM-generated text, making it crucial for mitigating potential misuse of LLMs. Currently, mainstream watermarking methods involve intervening in the inference phase of LLMs by modifying the logits output or influencing the token sampling process to produce watermarked text (Liu et al., 2023b; Pan et al., 2024; Liu et al., 2024b).
The KGW method (Kirchenbauer et al., 2023) modifies logits by categorizing the vocabulary into green and red lists, applying a bias towards green tokens during text generation. The bias value is typically determined by the parameter $\delta$, which reflects the watermark strength. For detection, it first labels each token as red or green and calculates the z-score of the entire text to reflect the proportion of green tokens; if this proportion exceeds a specified threshold, the text is classified as watermarked. In subsequent work, numerous enhancements have been suggested to improve text quality (Hu et al., 2024; Wu et al., 2023; Takezawa et al., 2023), increase information capacity (Wang et al., 2024; Yoo et al., 2023; Fernandez et al., 2023), boost robustness against watermark removal attacks (Zhao et al., 2024; Liu et al., 2024a; Ren et al., 2024; He et al., 2024), adapt to low-entropy conditions (Lee et al., 2023; Lu et al., 2024), and facilitate public detection (Liu et al., 2023a; Fairoze et al., 2023).
@@ 6.1 Experiment Settings
Hyper-parameters: The parameters related to WaterSeeker are as follows: $W=50,k=20$, with a tolerance for fragment connection set to 100. The parameter $\theta_{1}$ is set to 0.5, and $\theta_{2}$ is set to 1.5 for both algorithms. For full-text detection in baselines and during local traversal in WaterSeeker, the specific threshold selection is detailed in Appendix C. Notably, careful threshold selection is crucial for maintaining an acceptable false positive rate, as traversing long texts is prone to accumulating false positives. Evaluation-related parameters: $\theta$ mentioned in Section 3.2 is set to 0.5.
@@ 6.2 Detection Performance Analysis
Llama-2-7b Full-text Detection 0.017 1.000 0.000 0.001 0.017 1.000 0.000 0.001
WinMax 0.033 0.257 0.837 0.671 0.010 0.303 0.816 0.629
FLSW-100 0.017 0.480 0.677 0.434 0.000 0.473 0.690 0.430
FLSW-200 0.010 0.413 0.734 0.452 0.003 0.543 0.626 0.371
FLSW-300 0.010 0.563 0.604 0.359 0.003 0.787 0.351 0.294
FLSW-400 0.007 0.807 0.322 0.272 0.007 0.893 0.192 0.222
WaterSeeker(Ours) 0.030 0.270 0.830 0.653 0.007 0.330 0.799 0.602
Mistral-7b Full-text Detection 0.030 1.000 0.000 0.001 0.017 1.000 0.000 0.002
WinMax 0.060 0.147 0.892 0.745 0.043 0.453 0.688 0.488
FLSW-100 0.010 0.397 0.748 0.509 0.013 0.577 0.589 0.332
FLSW-200 0.013 0.343 0.786 0.510 0.017 0.657 0.505 0.295
FLSW-300 0.023 0.573 0.589 0.389 0.017 0.807 0.320 0.233
FLSW-400 0.027 0.840 0.271 0.289 0.017 0.933 0.123 0.171
WaterSeeker(Ours) 0.047 0.173 0.883 0.709 0.037 0.483 0.665 0.457
In Table 1, we present the detection performance of WaterSeeker alongside various baseline algorithms, evaluated on four datasets. The results demonstrate that WaterSeeker achieves lower FNR and higher F1 score, significantly surpassing other watermark detection algorithms, including Full-text Detection, FLSW-100, FLSW-200, FLSW-300, and FLSW-400, while being comparable to WinMax. Based on the principle that WinMax is guaranteed to reach the gold index by evaluating all possible windows, it represents the upper bound of detection performance. However, subsequent experiments show that WinMax has extremely high time complexity (Section 6.4). Furthermore, the IoU results in the table reveal that WaterSeeker possesses a significantly higher localization capability compared to the baselines, again comparable to WinMax.
@@ 6.4 Time Complexity Analysis
Table 2 lists the average time spent per sample for WaterSeeker and other baseline algorithms during detection. The results show that the time taken by WaterSeeker is comparable to that of full-text detection and FLSW, and significantly lower than the time required by WinMax. Sections 4.2 and 5.4 analyze the theoretical time complexity of WinMax and WaterSeeker, respectively. This section will further delve into a more in-depth analysis based on experimental results.
@@ 6.5 Ablation Study
Llama-2-7b w. local traverse 0.030 0.270 0.830 0.653 0.007 0.330 0.799 0.602
w/o local traverse 0.010 0.333 0.795 0.565 0.007 0.400 0.747 0.509
Mistral-7b w. local traverse 0.047 0.173 0.883 0.709 0.037 0.483 0.665 0.457
w/o local traverse 0.030 0.267 0.832 0.619 0.027 0.510 0.646 0.400
@@ Appendix B Detailed Proof for Aar
For the Aar method, we aim to analyze how the expected p-value for watermarked text varies with changes in the detection window size $W$.
@@
In the Aar method, the test statistic $S$ follows a Gamma distribution with shape parameter $W$ and scale parameter 1. For non-watermarked text, we have: $S\sim\text{Gamma}(W,1)$. The p-value is calculated using the Gamma CDF:
@@
Appendix C Detail of Threshold Selection in Full-text Detection
A key role of threshold selection is to control the false positive rate. In this context, the task involves detecting watermark fragments within long texts, which requires traversing extensive content and can lead to an accumulation of false positives. Therefore, managing the false positive rate within the detection window is crucial in this scenario.
Assuming the false positive rate of event $X$ is $\alpha$, let event $Y$ represent ”event $X$ has at least one false positive in 10,000 independent trials. ” We know that $P(Y)=1-(1-\alpha)^{10000}$. If we want to control $P(Y)$ to be around 0.01, then $\alpha$ should be set to approximately $10^{-6}$. Next, we will explain how to set thresholds for both KGW and Aar to ensure the theoretical false positive rate in window detection is below $10^{-6}$.

### v2

@@ Abstract
Watermarking algorithms for large language models (LLMs) have attained high accuracy in detecting LLM-generated text. However, existing methods primarily focus on distinguishing fully watermarked text from non-watermarked text, overlooking real-world scenarios where LLMs generate only small sections within large documents. In this scenario, balancing time complexity and detection performance poses significant challenges. This paper presents WaterSeeker, a novel approach to efficiently detect and locate watermarked segments amid extensive natural text. It first applies an efficient anomaly extraction method to preliminarily locate suspicious watermarked regions. Following this, it conducts a local traversal and performs full-text detection for more precise verification. Theoretical analysis and experimental results demonstrate that WaterSeeker achieves a superior balance between detection accuracy and computational efficiency. Moreover, WaterSeeker’s localization ability supports the development of interpretable AI detection systems. This work pioneers a new direction in watermarked segment detection, facilitating more reliable AI-generated content identification. Our code is available at https://github.com/THU-BPM/WaterSeeker.
@@ 1 Introduction
As the quality of text generated by large language models (LLMs) improves, these models address various practical challenges while also raising concerns such as misinformation (Liu et al., 2023b; Chen & Shu, 2024) and copyright infringement (Rillig et al., 2023). To tackle these issues, LLM watermarking technology has emerged, embedding specific information (watermarks) during text generation for accurate detection through specialized algorithms. Existing watermarking methods mainly focus on distinguishing fully watermarked text from non-watermarked content (Kirchenbauer et al., 2023; Zhao et al., 2024; Liu et al., 2024b; Lu et al., 2024; Aaronson & Kirchner, 2022; Kuditipudi et al., 2024). However, in real-world scenarios, LLMs often generate only brief segments within longer documents, making it essential to detect these watermarked sections among large amounts of non-watermarked text.
Most previous algorithms relied on computing statistics across the entire document for detection, which we refer to as full-text detection methods. For instance, some approaches categorize vocabulary into green and red lists, applying a bias towards the selection of green tokens during generation and using metrics like the z-score to measure their proportion throughout the entire text (Kirchenbauer et al., 2023; Zhao et al., 2024; Liu et al., 2024b). Other methods utilize pseudo-random numbers to guide token sampling, calculating correlations across the entire document during detection (Aaronson & Kirchner, 2022; Christ et al., 2024; Kuditipudi et al., 2024). Ultimately, the presence of a watermark is determined by whether these statistics exceed a specified threshold.
@@ 2 Related Work
Watermarking for large language models (LLMs) embeds subtle, algorithmically detectable signals in outputs to identify LLM-generated text, making it crucial for mitigating potential misuse of LLMs. Currently, mainstream watermarking methods involve intervening in the inference phase of LLMs by modifying the logits output or influencing the token sampling process to produce watermarked text (Liu et al., 2023a; Pan et al., 2024; Liu et al., 2024c).
The KGW method (Kirchenbauer et al., 2023) modifies logits by categorizing the vocabulary into green and red lists, applying a bias towards green tokens during text generation. The bias value is typically determined by the parameter $\delta$, which reflects the watermark strength. For detection, it first labels each token as red or green and calculates the z-score of the entire text to reflect the proportion of green tokens; if this proportion exceeds a specified threshold, the text is classified as watermarked. In subsequent work, numerous enhancements have been suggested to improve text quality (Hu et al., 2024; Wu et al., 2023; Takezawa et al., 2023), increase information capacity (Wang et al., 2024; Yoo et al., 2023; Fernandez et al., 2023), boost robustness against watermark removal attacks (Zhao et al., 2024; Liu et al., 2024b; Ren et al., 2024; He et al., 2024), adapt to low-entropy conditions (Lee et al., 2023; Lu et al., 2024), and facilitate public detection (Liu et al., 2024a; Fairoze et al., 2023).
@@ 6.1 Experiment Settings
Hyper-parameters: The parameters related to WaterSeeker are as follows: $W=50,k=20$, with a tolerance for fragment connection set to 100. The parameter $\theta_{1}$ is set to 0.5, and $\theta_{2}$ is set to 1.5 for both algorithms. The threshold selection within the specified window is detailed in Appendix C. Notably, careful threshold selection is crucial for maintaining an acceptable false positive rate, as traversing long texts is prone to accumulating false positives. Evaluation-related parameters: $\theta$ mentioned in Section 3.2 is set to 0.5.
@@ 6.2 Detection Performance Analysis
Llama-2-7b Full-text Detection 0.000 1.000 [0.987] 0.000 [0.026] 0.000 0.000 1.000 [0.967] 0.000 [0.065] 0.001
WinMax 0.017 0.273 0.834 0.661 0.017 0.410 0.734 0.537
FLSW-100 0.003 0.473 0.688 0.448 0.003 0.547 0.622 0.363
FLSW-200 0.003 0.440 0.716 0.417 0.000 0.577 0.595 0.338
FLSW-300 0.007 0.683 0.479 0.313 0.000 0.783 0.356 0.265
FLSW-400 0.003 0.897 0.187 0.230 0.003 0.933 0.125 0.191
WaterSeeker(Ours) 0.017 0.303 0.813 0.629 0.010 0.433 0.719 0.510
Mistral-7b Full-text Detection 0.000 1.000 [0.990] 0.000 [0.020] 0.000 0.000 1.000 [0.973] 0.000 [0.052] 0.001
WinMax 0.010 0.277 0.835 0.656 0.013 0.370 0.767 0.562
FLSW-100 0.000 0.500 0.667 0.433 0.000 0.497 0.670 0.411
FLSW-200 0.000 0.410 0.742 0.452 0.003 0.557 0.613 0.360
FLSW-300 0.003 0.573 0.597 0.342 0.003 0.827 0.295 0.254
FLSW-400 0.003 0.810 0.318 0.244 0.007 0.943 0.107 0.195
WaterSeeker(Ours) 0.007 0.287 0.829 0.642 0.010 0.390 0.753 0.542
In Table 1, we present the detection performance of WaterSeeker alongside various baseline algorithms, evaluated on four datasets. Given that full-text detection lacks localization capabilities, its FNR is necessarily 1 under the evaluation rule requiring IoU $>\theta$ for successful detection. Therefore, we additionally report the FNR and corresponding F1 score for Full-text Detection without the IoU $>\theta$ constraint, highlighted in red.
The results demonstrate that WaterSeeker achieves lower FNR and higher F1 score, significantly surpassing other watermark detection algorithms, including Full-text Detection, FLSW-100, FLSW-200, FLSW-300, and FLSW-400, while being comparable to WinMax. Based on the principle that WinMax is guaranteed to reach the gold index by evaluating all possible windows, it represents the upper bound of detection performance. However, subsequent experiments show that WinMax has extremely high time complexity (Section 6.4). Furthermore, the IoU [^1] For samples with no detected segments, the IoU is 0. For samples with multiple detected segments, take the max(IoU). results in the table reveal that WaterSeeker possesses a significantly higher localization capability compared to the baselines, again comparable to WinMax.
@@ 6.4 Time Complexity Analysis
Table 2 lists the average time spent per sample for WaterSeeker and other baseline algorithms during detection. The results show that the time taken by WaterSeeker is comparable to that of full-text detection and FLSW, and significantly lower than the time required by WinMax. Sections 4.2 and 5.4 analyze the theoretical time complexity of WinMax and WaterSeeker, respectively. This section aims to offer a more comprehensive analysis based on experimental results, taking into account the constant factors preceding the complexity terms. Additionally, it will investigate the reasons behind the significant disparity in execution time of WinMax between KGW and Aar.
@@ 6.5 Ablation Study
Llama-2-7b w. local traverse 0.017 0.303 0.813 0.629 0.010 0.433 0.719 0.510
w/o local traverse 0.003 0.390 0.756 0.520 0.000 0.450 0.710 0.462
Mistral-7b w. local traverse 0.007 0.287 0.829 0.642 0.010 0.390 0.753 0.542
w/o local traverse 0.003 0.380 0.763 0.534 0.003 0.413 0.738 0.490
@@ Appendix B Detailed Proof for Aar
B.1 Expected p-value of watermarked text
In this sub-section, we aim to analyze how the expected p-value for watermarked text varies with changes in the detection window size $W$.
@@ B.1 Expected p-value of watermarked text
B.2 Calculation of p-threshold to control false positive rate
@@
For non-watermarked text, $u_{i}\sim\text{Uniform}([0,1])$. Consequently, the test statistic $S$ follows a Gamma distribution: $S\sim\text{Gamma}(W,1)$, where $W$ is the shape parameter and 1 is the scale parameter. The p-value is calculated using the Gamma CDF:
@@
Appendix C Detail of Threshold Selection Within the Specified Window
A key role of threshold selection is to control the false positive rate. In this context, the task involves detecting watermark fragments within long texts, which requires traversing extensive content and can lead to an accumulation of false positives. Therefore, managing the false positive rate within the detection window is crucial in this scenario. In the experiment, we set the targeted false positive rate $\alpha$ within the detection window to $10^{-6}$.
C.1 Rationale for Setting $\alpha$ to $10^{-6}$
Table 5: Simulated FPR of WaterSeeker using 10,000 samples for each watermarking method. The targeted false positive rate within the detection window is set to $10^{-6}$.
Watermarking Method Simulated FPR
KGW 0.0054
Aar 0.0042
WaterSeeker, WinMax, and FLSW all involve employing sliding windows for text traversal and conduct full-text detection within each window. As these windows overlap, they cannot be treated as independent, making it challenging to derive a theoretical upper bound for the document-level FPR from the targeted FPR within each window. Given this, we utilize large-scale data simulation to demonstrate that, with a targeted false positive rate of $10^{-6}$ within each window, our proposed method WaterSeeker maintains an acceptable false positive rate.
For the KGW method, we set $\gamma=0.5$ in our experiments, meaning each token in non-watermarked text has a 0.5 probability of being green and 0.5 probability of being red. In the simulation, we generate 10,000 samples, each containing 10,000 tokens, with each token having a 0.5 probability of being 1 and 0.5 probability of being 0. For the Aar method, each token in non-watermarked text corresponds to $u_{i}\sim\text{Uniform}[0,1]$. In the simulation, we again generate 10,000 samples, each containing 10,000 tokens, with each token randomly assigned a floating-point number from $[0,1]$.
We then apply WaterSeeker to detect watermarked segments within these samples, setting the targeted false positive rate within the detection window to $10^{-6}$. The large-scale simulation results in Table 5 demonstrate that WaterSeeker maintains a false positive rate of approximately 0.005, which is considered acceptable. For scenarios requiring more stringent FPR control, the targeted false positive rate can be adjusted downward. However, this inevitably compromises the detection rate, highlighting a key challenge in watermarked segment detection within large documents.
C.2 Setting the Threshold to Achieve a Target False Positive Rate $\alpha$